# Iband Measurements in a Sample of Hosts of 10 SN Ia Match HST Cepheids

### v2

@@ 1 Introduction
Independent distance measurements to the same galaxies using multiple standard candles offers one of the best ways to ensure systematics are well accounted for and understood. The tip of the red giant branch (TRGB) refers to the maximum luminosity reached by first-ascent red giant stars before transitioning onto the horizontal branch, and has been used as a standard candle to measure distances (e.g. Anand et al., 2019a, b; Beaton et al., 2019; Hoyt et al., 2021; Lee et al., 2021a; Anand et al., 2021a, b; Oakes et al., 2022; Tran et al., 2022; Lee et al., 2022; Madore & Freedman, 2024; Anand et al., 2024a, b), and the Hubble constant (e.g. Freedman et al., 2019; Kim et al., 2020; Freedman, 2021a; Blakeslee et al., 2021; Dhawan et al., 2022; Anand et al., 2022; Dhawan et al., 2023; Scolnic et al., 2023, see also review in Valentino & Brout 2024). The TRGB can thus be used as a means to crosscheck systematics of other standard candles, such as Cepheid variables.
@@ 3 Photometry
We follow the general procedures outlined by previous works from our team for photometric reductions (Riess et al., 2024; Anand et al., 2024a). In detail, we perform PSF photometry333The photometry for this study is available publicly upon publication in a Zenodo repository, DOI: 10.5281/zenodo.13131990. using the DOLPHOT444http://americano.dolphinsim.com/dolphot/ software package (Dolphin, 2000, 2016), together with the latest (February 2024) version of the JWST/NIRCam module (Weisz et al., 2024). We use the stage 3 F150W i2d files as reference frames and perform photometry directly on the stage 2 cal images. We use the Vega-Vega system to remain consistent with earlier versions of DOLPHOT’s JWST module, which gives a 0.04 mag offset (in F090W) from the Vega-Sirius system that is currently default in DOLPHOT NIRCam module. We use the “-etctime” option, which allows the package to use (TMEASURE) for S/N calculations. We apply the following DOLPHOT quality cuts, similar to those from Weisz et al. (2024): crowd 0.5 (in both bands), sharp2 0.01 (in both bands), and type 2. In addition to these quality cuts, we also apply S/N cuts in both bands for each target. For NGC 1448, NGC 1559, NGC 3447, NGC 5584, and NGC 5643, we adopt a minimum threshold of S/N=5, whereas we choose a value of S/N=3 for NGC 2525, NGC 3370, and NGC 5861 (due to the decreased relative depth of their photometry).
We also correct for foreground extinction using E(B-V) from Schlafly & Finkbeiner (2011) and galaxy coordinates, both retrieved using the the NASA/IPAC Extragalactic Distance Database (NED)555https://ned.ipac.caltech.edu/. We use the Fitzpatrick (1999) reddening law with = 3.1 and E(B-V) = 1.4156 and 0.6021 for F090W and F150W, respectively (Anand et al., 2024a). We adopt the extinction used in Anand et al. (2022) of E(B-V) = 0.161 mag for NGC 5643, which is located at a low galactic latitude (where the dust maps are less certain), and where the adopted value was re-estimated using the displacement of the zero-age main sequence following Rizzi et al. (2017).
@@ 4 Measurement
Measuring the TRGB magnitude involves locating the discontinuity or inflection point of the luminosity function consisting of red giant and asymptotic giant branch stars. This can be done many ways; for instance, using a Sobel filter (Lee et al., 1993; Hatt et al., 2017), or model-based methods, with either least-squares fitters (Wu et al., 2014; Crnojević et al., 2019) or maximum-likelihood estimation (Makarov et al., 2006; Li et al., 2022, 2023a). In this study, we measure the TRGB using a Sobel filter using the measurement routine without spatial clipping (which we instead perform using elliptical annuli) from the Comparative Analysis of TRGBs (CATs) team (Wu et al., 2023; Scolnic et al., 2023; Li et al., 2023b), which is publicly available on GitHub666https://github.com/JiaxiWu1018/Unsupervised-TRGB. The CATs team had optimized their measurement parameters (such as smoothing value, color cuts, etc.) by minimizing field-to-field dispersion in the TRGB measurement across multiple fields. At the moment, there are a limited number of JWST fields in each host galaxy, which prevents us from performing a similar optimization. For this study, we adopt a smoothing value of 0.05 mag and unweighted, or ‘simple’ weighting as named in the CATs algorithm (i.e. no further weighting of the output of the edge-detection measurement). We note that without the CATs optimizations in place, this methodology is equivalent to the TRGB measurements adopted by the Carnegie-Chicago Hubble Program (CCHP; Beaton et al., 2016; Freedman et al., 2019), but without the additional weighting of the Sobel filter output (which has been shown to cause biases in the outputs; Anderson et al. 2024; Anand et al. 2024a). We explore the effects of varying the level of smoothing of the luminosity function and the color range used to measure the luminosity function, 14 combinations in all, in the next section. TRGB uncertainties are estimated using bootstrap resampling of the luminosity function with 100 samples. For NGC 5584, using the smaller smoothing value of =0.05 mag produces a Sobel response that shows two peaks of nearly the same height. Following Anand et al. (2024a), if there are two Sobel peaks that are of equivalent height ( 3), we select the brighter peak.
@@
Before measuring the TRGB, we perform spatial selections to limit contamination from younger stellar populations in these star-forming galaxies. The NIRCam regions used to measure the TRGB are highlighted in Fig. 1 in all cases (except NGC 3447, described below), only regions exterior to the mag/sq arcsec isophote (; de Vaucouleurs et al. 1991) are selected (Anand et al., 2022, 2024a; Hoyt et al., 2024). For NGC 1559 and NGC 5584, the parallel NIRCam module contains a suitable sampling of stars, although only half of the parallel module is used (the far half of each contain too few stars to perform precise empirical PSF adjustments and aperture corrections). The same is nearly true for NGC 5643, though we remove a small portion of each parallel module that is interior to . In NGC 1448, NGC 2525, NGC 3370, and NGC 5861, a sizeable portion of the main module is outside of , which we include to increase the sampling of stars and for NGC 2525, NGC 3370 and NGC 5861, the parallel modules contain too few stars to perform photometry on. Lastly, NGC 3447 is an actively interacting system, often described as two targets, N3447A and B. Here, we use two ellipses centered at RA, Dec = (10:53:23.998, +16:46:26.78) and (10:53:29:3934, +16:47:02.294), major axes of 140 and 50 arcsec, minor axes of 60 and 40 arcsec, and position angles of 80 and 115 degrees, respectively, following Mazzei et al. (2018). A color image of NGC 3370 is shown in Fig. 2, along with the same selection region adopted for the spatial cuts. It can be seen that the outer regions of the image used for our analysis are of low stellar density, where stars are easily separated.
@@ 5 Distances
| Galaxy | TRGB | TRGB | TRGB | (TRGB) | (TRGB) | (Cepheid) | (Cepheid) | (var) | ||||
| s=0.10 | s=0.10 | s=0.05 | s=0.05 | (s=0.10s=0.05) | s=0.10 | s=0.10 | s=0.05 | s=0.05 | ||||
| NGC 1448 | 26.97 | 0.02 | 27.02 | 0.05 | 0.05 | 31.34 | 0.05 | 31.39 | 0.07 | 31.30 | 0.05 | 0.02 |
| NGC 1559 | 27.13 | 0.01 | 27.14 | 0.01 | 0.01 | 31.50 | 0.05 | 31.51 | 0.05 | 31.50 | 0.07 | 0.01 |
| NGC 2525 | 27.50 | 0.03 | 27.44 | 0.07 | 0.06 | 31.87 | 0.06 | 31.81 | 0.08 | 32.06 | 0.11 | 0.04 |
| NGC 3370 | 27.82 | 0.03 | 27.82 | 0.07 | 0 | 32.19 | 0.06 | 32.19 | 0.08 | 32.13 | 0.06 | 0.02 |
| NGC 3447 | 27.56 | 0.04 | 27.55 | 0.06 | 0.01 | 31.93 | 0.07 | 31.92 | 0.08 | 31.95 | 0.05 | 0.01 |
| NGC 5584 | 27.58 | 0.06 | 27.43 | 0.10 | 0.15 | 31.95 | 0.08 | 31.8 | 0.11 | 31.77 | 0.06 | 0.07 |
| NGC 5643 | 26.20 | 0.01 | 26.21 | 0.01 | 0.01 | 30.57 | 0.06 | 30.58 | 0.06 | 30.55 | 0.06 | 0.01 |
| NGC 5861 | 27.73 | 0.03 | 27.73 | 0.11 | 0 | 32.10 | 0.06 | 32.10 | 0.12 | 32.23 | 0.11 | 0.02 |
@@
| 0.05 | 1.15, 1.65 | 0.022 | 0.033 | Anand et al. (2024a) | 1.130 | 0.10 |
| 0.05 | 1.15, 1.75 | 0.021 | 0.034 | Anand et al. (2024a) | 1.465 | 0.12 |
| 0.05 | 1.30, 1.75 | 0.005 | 0.036 | Anand et al. (2024a) | 0.741 | 0.09 |
| 0.07 | 1.15, 1.65 | 0.015 | 0.032 | Anand et al. (2024a) | 1.014 | 0.09 |
| 0.07 | 1.15, 1.75 | 0.013 | 0.033 | Anand et al. (2024a) | 1.37 | 0.11 |
| 0.07 | 1.30, 1.75 | 0.014 | 0.034 | Anand et al. (2024a) | 1.163 | 0.10 |
| 0.10 | 1.15, 1.65 | 0.006 | 0.031 | Anand et al. (2024a) | 1.003 | 0.09 |
| 0.10 | 1.15, 1.75 | 0.014 | 0.031 | Anand et al. (2024a) | 1.117 | 0.09 |
| 0.10 | 1.30, 1.75 | 0.014 | 0.032 | Anand et al. (2024a) | 0.960 | 0.09 |
| 0.10 | 1.30, 1.75 | 0.003 | 0.033 | Newman et al. (2024b) | 0.939 | 0.08 |
@@ 6 Discussion
Examining the Hubble Tension necessitates close scrutiny of the consistency and systematics of distance indicators. Based on the measurements made in this study, we find no evidence of a bias or systematic difference between the Cepheid and TRGB methods that would cause a difference translatable to that can solve the Hubble Tension. This finding is in agreement with Freedman & Madore (2023), who find, “the excellent agreement between the published Cepheid distances in Riess et al. (2022) and TRGB distances in Freedman et al. (2021), which in the mean, agree to 0.007 mag”. In addition, Fig. 5 shows a comparison between Cepheid and TRGB distances, both measured with HST, using TRGB and Cepheid distances from Freedman (2021b) and Riess et al. (2022), respectively. The TRGB and Cepheid distances in Fig. 5 yield a mean difference of 0.00 0.02 (stat) 0.04 (sys) mag, also consistent with no difference even without additional consideration of populations differences in TRGB measurements (Anderson et al., 2024; Koblischke & Anderson, 2024). This comparison will benefit from more JWST TRGB measurements in the future (for instance, from GO-1995; Freedman et al. 2021). As the Hubble Tension corresponds to 5(73/67.5) 0.17 mag, the demonstrated consistency is meaningful and is inconsistent with providing evidence for a solution to the Hubble Tension at 3.

### v3

@@ 1 Introduction
Independent distance measurements to the same galaxies using multiple standard candles offers one of the best ways to ensure systematics are well accounted for and understood. The tip of the red giant branch (TRGB) refers to the maximum luminosity reached by first-ascent red giant stars before transitioning onto the horizontal branch, and has been used as a standard candle to measure distances (e.g. Anand et al., 2019a, b; Beaton et al., 2019; Hoyt et al., 2021; Lee et al., 2021a; Anand et al., 2021a, b; Oakes et al., 2022; Tran et al., 2022; Lee et al., 2022; Madore & Freedman, 2024; Anand et al., 2024a, b), and the Hubble constant (e.g. Freedman et al., 2019; Kim et al., 2020; Freedman, 2021a; Blakeslee et al., 2021; Dhawan et al., 2022; Anand et al., 2022; Dhawan et al., 2023; Scolnic et al., 2023, see also review in Di Valentino & Brout 2024). The TRGB can thus be used as a means to crosscheck systematics of other standard candles, such as Cepheid variables.
@@ 3 Photometry
We follow the general procedures outlined by previous works from our team for photometric reductions (Riess et al., 2024a; Anand et al., 2024a). In detail, we perform PSF photometry333The photometry for this study is available publicly upon publication in a Zenodo repository, DOI: 10.5281/zenodo.13131990. using the DOLPHOT444http://americano.dolphinsim.com/dolphot/ software package (Dolphin, 2000, 2016), together with the latest (February 2024) version of the JWST/NIRCam module (Weisz et al., 2024). We use the stage 3 F150W i2d files as reference frames and perform photometry directly on the stage 2 cal images. We use the Vega-Vega system to remain consistent with earlier versions of DOLPHOT’s JWST module, which gives a 0.04 mag offset (in F090W) from the Vega-Sirius system that is currently default in DOLPHOT NIRCam module. We use the “-etctime” option, which allows the package to use (TMEASURE) for S/N calculations. We apply the following DOLPHOT quality cuts, similar to those from Weisz et al. (2024): crowd 0.5 (in both bands), sharp2 0.01 (in both bands), and type 2. In addition to these quality cuts, we also apply S/N cuts in both bands for each target. For NGC 1448, NGC 1559, NGC 3447, NGC 5584, and NGC 5643, we adopt a minimum threshold of S/N=5, whereas we choose a value of S/N=3 for NGC 2525, NGC 3370, and NGC 5861 (due to the decreased relative depth of their photometry).
We also correct for foreground extinction using E(B-V) from Schlafly & Finkbeiner (2011) and galaxy coordinates, both retrieved using the the NASA/IPAC Extragalactic Distance Database (NED)555https://ned.ipac.caltech.edu/and listed in Table 2 and the TRGB color selection has been adjusted to account for the reddening correction. We use the Fitzpatrick (1999) reddening law with = 3.1 and E(B-V) = 1.4156 and 0.6021 for F090W and F150W, respectively (Anand et al., 2024a). We adopt the extinction used in Anand et al. (2022) of E(B-V) = 0.161 mag for NGC 5643, which is located at a low galactic latitude (where the dust maps are less certain), and where the adopted value was re-estimated using the displacement of the zero-age main sequence following Rizzi et al. (2017).
@@ 4 Measurement
Measuring the TRGB magnitude involves locating the discontinuity or inflection point of the luminosity function consisting of red giant and asymptotic giant branch stars. This can be done many ways; for instance, using a Sobel filter (Lee et al., 1993; Hatt et al., 2017), or model-based methods, with either least-squares fitters (Wu et al., 2014; Crnojević et al., 2019) or maximum-likelihood estimation (Makarov et al., 2006; Li et al., 2022, 2023a). In this study, we measure the TRGB using a Sobel filter using the measurement routine without spatial clipping (which we instead perform using elliptical annuli) from the Comparative Analysis of TRGBs (CATs) team (Wu et al., 2023; Scolnic et al., 2023; Li et al., 2023b), which is publicly available on GitHub666https://github.com/JiaxiWu1018/Unsupervised-TRGB. The CATs team had optimized their measurement parameters (such as smoothing value, color cuts, etc.) by minimizing field-to-field dispersion in the TRGB measurement across multiple fields. At the moment, there are a limited number of JWST fields in each host galaxy, which prevents us from performing a similar optimization. For this study, we adopt a smoothing value of 0.05 mag and unweighted, or ‘simple’ weighting as named in the CATs algorithm (i.e. no further weighting of the output of the edge-detection measurement). We note that without the CATs optimizations in place, this methodology is equivalent to the TRGB measurements adopted by the Carnegie-Chicago Hubble Program (CCHP; Beaton et al., 2016; Freedman et al., 2019), but without the additional weighting of the Sobel filter output (which has been shown to cause biases in the outputs; Anderson et al. 2024; Anand et al. 2024a). We explore the effects of varying the level of smoothing of the luminosity function and the color range used to measure the luminosity function, 14 combinations in all, in the next section. TRGB uncertainties are estimated using bootstrap resampling of the luminosity function with 5,000 samples and the standard deviation of the bootstrap distribution. These distributions all follow Gaussianity with means that lie within 1 from the measured TRGB. For NGC 5584, using the smaller smoothing value of =0.05 mag produces a Sobel response that shows two peaks of nearly the same height. Following Anand et al. (2024a), if there are two Sobel peaks that are of equivalent height ( 3), we select the brighter peak. The bootstrap distribution for NGC 5584 exhibits two Gaussians corresponding to the two Sobel peaks in the baseline result. We incorporate this spread to account for uncertainty in peak selection for NGC 5584.
@@
Before measuring the TRGB, we perform spatial selections to limit contamination from younger stellar populations in these star-forming galaxies. The NIRCam regions used to measure the TRGB are highlighted in Fig. 1 in all cases (except NGC 3447, described below), only regions exterior to the mag/sq arcsec isophote (; de Vaucouleurs et al. 1991) are selected (Anand et al., 2022, 2024a; Hoyt et al., 2024). For NGC 1559 and NGC 5584, the parallel NIRCam module contains a suitable sampling of stars, although only half of the parallel module is used (the far half of each contain too few stars to perform precise empirical PSF adjustments and aperture corrections). The same is nearly true for NGC 5643, though we remove a small portion of each parallel module that is interior to . In NGC 1448, NGC 2525, NGC 3370, and NGC 5861, a sizeable portion of the main module is outside of , which we include to increase the sampling of stars and for NGC 2525, NGC 3370 and NGC 5861, the parallel modules contain too few stars to perform photometry on. Lastly, NGC 3447 is an actively interacting system, often described as two targets, N3447A and B. Here, we use two ellipses centered at RA, Dec = (10:53:23.998, +16:46:26.78) and (10:53:29:3934, +16:47:02.294), major axes of 140 and 50 arcsec, minor axes of 60 and 40 arcsec, and position angles of 80 and 115 degrees, respectively, following Mazzei et al. (2018).
A color image of NGC 3370 is shown in Fig. 2, along with the same selection region adopted for the spatial cuts. It can be seen that the outer regions of the image used for our analysis are of low stellar density, where stars are easily separated. While in some cases the TRGB is only 1 magnitude above the photometric completeness limit, we note that our spatial selection criteria as described above limits the effects of photometric bias, which would otherwise skew our edge-detection measurements in higher surface-brightness regions. Even for galaxies where the TRGB is located closest to the photometric limit, we do not measure any noticeable photometric bias in our selected halo regions–– for example, the TRGB in NGC 3370 lies near a photometric completeness of 65, but the measured photometric bias remains negligible (0.01 mag).
@@ 5 Distances
| Galaxy | E(B-V) | R | NBT | TRGB | TRGB | TRGB | (TRGB) | (TRGB) | (var) | ||||||
| s=0.10 | s=0.10 | s=0.05 | s=0.05 | (s=0.10s=0.05) | s=0.10 | s=0.10 | s=0.05 | s=0.05 | Cepheid | Cepheid | |||||
| NGC 1448 | 0.012 | 3.2 | 7392 | 26.97 | 0.02 | 27.02 | 0.05 | 0.05 | 31.34 | 0.05 | 31.39 | 0.07 | 31.30 | 0.05 | 0.02 |
| NGC 1559 | 0.026 | 3.8 | 1582 | 27.13 | 0.01 | 27.14 | 0.01 | 0.01 | 31.50 | 0.05 | 31.51 | 0.05 | 31.50 | 0.07 | 0.01 |
| NGC 2525 | 0.052 | 2.1 | 8225 | 27.50 | 0.03 | 27.44 | 0.07 | 0.06 | 31.87 | 0.06 | 31.81 | 0.09 | 32.06 | 0.11 | 0.04 |
| NGC 3370 | 0.028 | 1.7 | 617 | 27.82 | 0.03 | 27.82 | 0.07 | 0 | 32.19 | 0.06 | 32.19 | 0.08 | 32.13 | 0.06 | 0.02 |
| NGC 3447 | 0.026 | 1.9 | 678 | 27.56 | 0.04 | 27.55 | 0.08 | 0.01 | 31.93 | 0.06 | 31.92 | 0.09 | 31.95 | 0.05 | 0.01 |
| NGC 5584 | 0.035 | 2.6 | 1582 | 27.58 | 0.06 | 27.43 | 0.10 | 0.15 | 31.95 | 0.08 | 31.8 | 0.11 | 31.77 | 0.06 | 0.07 |
| NGC 5643 | 0.161 | 3.2 | 21032 | 26.20 | 0.01 | 26.21 | 0.01 | 0.01 | 30.57 | 0.06 | 30.58 | 0.06 | 30.55 | 0.06 | 0.01 |
| NGC 5861 | 0.095 | 1.5 | 6199 | 27.73 | 0.03 | 27.73 | 0.10 | 0 | 32.10 | 0.06 | 32.10 | 0.11 | 32.23 | 0.11 | 0.02 |
@@
| 0.05 | 1.15, 1.65 | 0.022 | 0.034 | Anand et al. (2024a) | 1.103 | 0.10 |
| 0.05 | 1.15, 1.75 | 0.018 | 0.034 | Anand et al. (2024a) | 1.363 | 0.11 |
| 0.05 | 1.30, 1.75 | 0.007 | 0.037 | Anand et al. (2024a) | 0.712 | 0.09 |
| 0.07 | 1.15, 1.65 | 0.014 | 0.033 | Anand et al. (2024a) | 1.008 | 0.09 |
| 0.07 | 1.15, 1.75 | 0.014 | 0.033 | Anand et al. (2024a) | 1.317 | 0.11 |
| 0.07 | 1.30, 1.75 | 0.013 | 0.034 | Anand et al. (2024a) | 1.083 | 0.10 |
| 0.10 | 1.15, 1.65 | 0.011 | 0.031 | Anand et al. (2024a) | 1.057 | 0.09 |
| 0.10 | 1.15, 1.75 | 0.015 | 0.031 | Anand et al. (2024a) | 1.121 | 0.09 |
| 0.10 | 1.30, 1.75 | 0.014 | 0.032 | Anand et al. (2024a) | 0.961 | 0.09 |
| 0.10 | 1.30, 1.75 | 0.005 | 0.032 | Newman et al. (2024b) | 0.882 | 0.08 |
@@ 6 Discussion
Examining the Hubble Tension necessitates close scrutiny of the consistency and systematics of distance indicators. Based on the measurements made in this study, we find no evidence of a bias or systematic difference between the Cepheid and TRGB methods that would cause a difference translatable to that can solve the Hubble Tension. This finding is in agreement with Freedman & Madore (2023), who find, “the excellent agreement between the published Cepheid distances in Riess et al. (2022) and TRGB distances in Freedman et al. (2021), which in the mean, agree to 0.007 mag”. While the recent analysis in Freedman et al. (2024) found a 2.5 disagreement between TRGB and Cepheid distances, they do not include systematic measurement errors in NGC 4258 for the two standard candles of 0.04 and 0.09 mag, respectively. Sufficiently accounting for these terms in their comparison decreases the difference to below 1 and provides no evidence of a significant systematic difference (see Appendix B in Riess et al. (2024b). Notably, the primary paper describing the TRGB results in Freedman et al. (2024) (Hoyt et al., in prep) is not yet available, preventing us from making any further comparisons. In addition, Fig. 5 shows a comparison between Cepheid and TRGB distances, both measured with HST, using TRGB and Cepheid distances from Freedman (2021b) and Riess et al. (2022), respectively. The TRGB and Cepheid distances in Fig. 5 yield a mean difference of 0.00 0.02 (stat) 0.04 (sys) mag, also consistent with no difference even without additional consideration of populations differences in TRGB measurements (Anderson et al., 2024; Koblischke & Anderson, 2024). This comparison will benefit from more JWST TRGB measurements in the future (for instance, from GO-1995; Freedman et al. 2021). As the Hubble Tension corresponds to 5(73/67.5) 0.17 mag, the demonstrated consistency is meaningful and is inconsistent with providing evidence for a solution to the Hubble Tension at 3.